bill is 6 years older than Tom. Three years ago Bill's age was four times Tom's age. What is the age of each boy?
Answer by algebrahouse.com(1659)

"bill is 6 years older than Tom. Three years ago Bill's age was four times Tom's age. What is the age of each boy?"

x = Tom's age now
x + 6 = Bill's age now

x - 3 = Tom's age 3 years ago {subtracted 3 from present age}
x + 3 = Bill's age 3 years ago {subtracted 3 from present age}

x + 3 = 4(x - 3) {three years ago Bill's age was four times Tom's age}
x + 3 = 4x - 12 {used distributive property}
3 = 3x - 12 {subtracted x from both sides}
3x = 15 {added 12 to both sides}
x = 5 {divided both sides by 3}
x + 6 = 11 {substituted 5, in for x, into x + 6}

Tom is 5
Bill is 11
www.algebrahouse.com